OpenStudy (anonymous):

What is the solution set of |2x – 3| – 8 = –1?

OpenStudy (anonymous):

Add 8 to both sides: |2x-3| = 7. This is equivalent to saying that -(2x-3)=7 or 2x-3 = 7. Any x that satisfy one of these, will satisfy the original expression with the absolute value sign.

OpenStudy (anonymous):

x = -2, or x = 5. They are the only two solutions.

OpenStudy (anonymous):

Case I and Case II 2x - 3 >0 2x - 3 <0